AMSTEX(1) AMSTEX(1) NAME amstex−structuredtextformattingandtypesetting SYNOPSIS amstex[firstline] DESCRIPTION Thismanualpageisnotmeanttobeexhaustive.Thecompletedocumentationforthisversionof TEXcanbefoundintheinfofileormanualWeb2C:ATeXimplementation. TheAMSTEXlanguageisdescribedinthebookTheJoyofTEX.AMSTEXisaTEXmacropackage,notamodificationtotheTEXsourceprogram,soallthecapabilitiesdescribedintex(1)are present. TheAMSTEXmacrosencouragewriterstothinkaboutthecontentoftheirdocuments,rather thantheform.Theideal,notalwaysrealized,istohavenoformattingcommands(like‘‘switchto italic’’or‘‘skip2picas’’)inthedocumentatall;instead,everythingisdonebyspecificmarkup instructions:‘‘emphasize’’,‘‘startasection’’. AMSTEXistheofficialtypesettingsystemoftheAmericanMathematicalSociety,andnearlyallof itspublicationsaretypesetusingAMSTEX.TheonlineversionofAMSMathReviewsuses AMSTEXinputsyntaxfordisplayofmathematicalmaterial.AMSauthorscanprovideeditors withcomputer-readableAMSTEXfiles. ForauthorsmorefamiliarwithLATEX,anAMS-supportedpackagecalledamsmath.styisisavail-able. FILES amsguide.tex" DocumentationforAMSTEX,typesettablebyTEX. amfndoc.tex" DocumentationforAMSfonts,typesettablebyTEX. SEEALSO latex(1),tex(1),slitex(1). MichaelSpivak,TheJoyofTEX,2ndedition,Addison-Wesley,1990,ISBN0-8218-2997-1. Web2C7.32x 27December1992 1